OpenStudy (anonymous):

Solve -4(x + 10) - 6 = -3(x - 2).

pooja195 (pooja195):

-4(x + 10) distribute

OpenStudy (nurali):

-4(x + 10) - 6 = -3(x - 2) -4x-40-6=-3x+6 -4x+3x=6+46 -x=52 x=-52

OpenStudy (anonymous):

i got -4x-40 -6=-3x+6 ....

OpenStudy (anonymous):

lol yeah tanks thats all i needed to know! thanks

OpenStudy (nurali):

My Pleasure.